OpenStudy (anonymous):

can someone help me solve this? 15√2 ÷ √2

OpenStudy (anonymous):

whaaaat? where did the a come from?

OpenStudy (anonymous):

so what would it be?

OpenStudy (anonymous):

can u show me how to do it with the numbers that i need?

OpenStudy (anonymous):

another way of writing it:\[{15{\sqrt{2}}}\over{\sqrt{2}}\]so you know that anything divided by itself can be cancelled out:\[{15{\cancel{\sqrt{2}}}}\over{\cancel{\sqrt{2}}}\]and all you are left with is:\[15\over1\]which is just:\[\large15\]

OpenStudy (anonymous):

THANK U THANK U THANK U!!:D